A point charge q1=8.60 nC is fixed at the origin. A second point charge q2=5.30 nC is fixed at the point x=10.0 cm.

(a) Find the electric potential at point A, located at xA=14.0 cm.

(b) Find the total electric field (magnitude and direction) at point B, located at xB=6.00 cm.

Please include your work and equations.
Answer #1

A)

V = k ( q1/r1 + q2/r2)

V = 9* 10^9 ( 8.6 / 0.14 + 5.3 /0.04)* 10^-9

V = 1745.36 V

=====

B)

E = k ( q/r1^2 + q/r2^2)

E = 9* 10^9 * 10^-9* ( 8.6 / 0.14^2 + 5.3 / 0.04^2)

E = 33761.48 N/C

Comment in case any doubt please rate my answer....

​​​​​​...